OpenStudy (anonymous):

what are the real or imaginary soultions of each polynomial equation x^4-20x^2=-64

OpenStudy (whpalmer4):

\[x^4-20x^2=-64\]\[x^4-20x^2+64=0\]Do you know how to factor that equation? One way to do so is to substitute \(u = x^2\), giving you \[u^2-20u+64=0\]which you should be able to factor. Then undo your previous substitution and solve the two equations. I'll do an example: \[x^4-2x^2+1=0\]Substitute \(u=x^2\)\[u^2-2u+1=0\]Factor\[(u-1)(u-1)=0\]Solve equations\[u-1=0\]\[u=1\]Undo substitution\[x^2=1\]Solve\[x=\pm1\]As both factors were identical, the solutions to the original were \(x=\pm1\), with both having multiplicity 2. As a quick check: \[(1)^4-2(1)^2+1=0\]\[1-2+1=0\]\[0=0\checkmark\]\[(-1)^4-2(-1)^2+1=0\]\[1-2+1=0\]\[0=0\checkmark\]